OpenStudy (anonymous):

Who wants a medal? How much of a 60% acid solution must be mixed with 50 gallons of a 24% acid solution to obtain a solution that is 50% acid? a. 130 gal c. 13 gal b. 125 gal d. None of these

OpenStudy (joannablackwelder):

I would use the idea of a mass balance of the acid.

OpenStudy (joannablackwelder):

For example, the amount of acid in the 50 gallons of a 24% acid solution would be 50*.24=12 gallons of acid.

OpenStudy (joannablackwelder):

Make sense so far?

OpenStudy (anonymous):

yes

OpenStudy (joannablackwelder):

Sweet. Since we don't have the volume of the 60% acid solution, I will call that x.

OpenStudy (joannablackwelder):

Can you write an expression that shows the volume of acid of the 60% acid solution?

OpenStudy (anonymous):

0.5 = 0.6x + (50*0.24) I have a feeling that this isn't right.

OpenStudy (joannablackwelder):

That is close. The addition side looks good, but what about the volume of the combined solution?

OpenStudy (joannablackwelder):

@Catseyeglint911?

OpenStudy (anonymous):

sorry I had to step a way for a bit

OpenStudy (joannablackwelder):

No worries :)

OpenStudy (anonymous):

(0.5*50) = 0.6x + (50*0.24) Is this what you meant?

OpenStudy (joannablackwelder):

Closer. :) The volume of the 24% solution is 50 gals and the 60% acid solution has volume x gals, so wouldn't the combined 50% acid solution have more volume that 50 gals?

OpenStudy (anonymous):

So we would have to use two variables since we don't know the number of gallons need of the 60% solution nor do we know the number of gallons of the final solution?

OpenStudy (joannablackwelder):

We could, but we don't have to. Since we know we are adding two volumes together and one volume is 50 gallons and we defined the second volume to be x gallons. So the combined volume is 50+x gallons.

OpenStudy (anonymous):

(x+50) = 0.6 + (50*0.24) It's still wrong isn't it?

OpenStudy (joannablackwelder):

Last time you put (0.5*50) = 0.6x + (50*0.24) which was closer. The only thing you needed to change was (0.5*[50+x]) = 0.6x + (50*0.24)

OpenStudy (joannablackwelder):

The 50 to a 50+x

OpenStudy (joannablackwelder):

Do you follow me?

OpenStudy (anonymous):

Ooohhh! I see now. :D

OpenStudy (joannablackwelder):

:D

OpenStudy (joannablackwelder):

Can you solve for x?

OpenStudy (anonymous):

(0.5*[50+x]) = 0.6x + (50*0.24) 25 + 0.5x = 0.6x + 12 13 + 0.5x = 0.6x 13 = 0.1x 130 = x right?

OpenStudy (joannablackwelder):

Awesome! :)

OpenStudy (anonymous):

Thank you so much. :)

OpenStudy (joannablackwelder):

No worries